OpenStudy (anonymous):

f(x) = twenty-four over the quantity of 3 x minus two. Find f(-2).

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) finding f(-2) means that you plug in (-2) for x so 24 / (3(-2) - 2)) = ?

OpenStudy (one098):

f(-2)=24/3(-2)-2

OpenStudy (anonymous):

x−223=y you got from y=243x−2 remember you can only subtract and add fractions when they have the SAME denominator. Here, you have to treat (3x-2) as one term.